Grade 12General Physics

If the kinetic energy of a body charges by 20 % then its momentum would changed by

Shasanka Shekhar Mridha
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student 
 

Kinetic energy is given as KE=p^2/2 m=(1/2 m) P^2= constant.P^2

Now P—->P+0.20—→KE—→const.(P+0.20 P)^2

So( P+0.20 P)*2 -P^2= 0. 44 P^2 ,the increase in KE.

Therefore,0.44 P^2/P^2= 0.44 = 44%. the percentage increase in the KE.

Regards

Arun